OpenStudy (anonymous):

Simplify:

OpenStudy (anonymous):

\[\frac{ 12f^6g }{ 9f^3 }\]

OpenStudy (anonymous):

That is equal to \[\frac{ 12f^6gf^{-3} }{ 9 }\] so, 12 and 9 have a common divisor and f^ a* f^-b =f^(a-b) so your result will be...

OpenStudy (anonymous):

\[\frac{ 4f^3 g }{ 3 }\] I think its this.

OpenStudy (anonymous):

good job c: